OpenStudy (anonymous):

A system Of Linear Equations is shown below 4x+3y =8 y=2x+3 What is the value of y in the solution? A. -1/10 B. -16/5 C.14/5 D.1/10

OpenStudy (anonymous):

@suckerofmath is it B ?

OpenStudy (anonymous):

@campbell_st

OpenStudy (anonymous):

Sub B in your Y in the first equation and solve. Y=4x+3(-16/5)=8

OpenStudy (anonymous):

It's not b ?

OpenStudy (anonymous):

I don't know. Give me a minute to solve

OpenStudy (anonymous):

Oh ty sorry

OpenStudy (anonymous):

I would say B. It seams to work

OpenStudy (anonymous):

ok ty I have 6 more to go and I'm done with all my Easter vacation HW finally

OpenStudy (johnweldon1993):

Not quite...

OpenStudy (anonymous):

Wait nope I read the question wrong it's A

OpenStudy (anonymous):

Sorry

OpenStudy (johnweldon1993):

Not quite correct again...

OpenStudy (johnweldon1993):

4x + 3y = 8 y = 2x + 3 We want to solve for 'y' ...so we need to solve that 1st equation for 'x'...and substitute that into the second equation ... 4x + 3y = 8 4x = 8 - 3y x = (8 - 3y)/4 x = 2 - (3/4)y Now substitute that into the second equation for 'x' y = 2x + 3 becomes y = 2(2 - (3/4)y) + 3 y = 4 - (6/4)y + 3 y = 7 - (3/2)y y + (3/2)y = 7 (5/2)y = 7 5y = 14 y = 14/5

OpenStudy (anonymous):

HUh

OpenStudy (anonymous):

C wow we were wrong ...TY John. I'm John too

OpenStudy (johnweldon1993):

Just make sure never to accept answers straight out....try and work them out yourself